OpenStudy (anonymous):

What is the slope of a line parallel to the line whose equation is y - x = 5? -1 1/5 1 5

OpenStudy (gorv):

parallel line have same slope given line y=x+5 slope =1 so slope of parallel line =1

OpenStudy (anonymous):

y=x+5

OpenStudy (anonymous):

so 1